OpenStudy (anonymous):

i have a question how do you get 5/8 from this problem? 1/8+ 1/2, can you you please break down AND tell me?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

multiply 1/2 by 4/4. 1/2*4/4=4/8. Since 4/4=1 this doesn't change anything. 1/8+4/8=5/8 (same denominator so we can just add the numerators)

OpenStudy (anonymous):

thank you good answer!!!!!